# GateRA: Token-Aware Modulation for Parameter-Efficient Fine-Tuning

### v2

@@ Introduction
Large pre-trained models have revolutionized natural language processing and vision-language tasks, yet their massive parameter counts make full fine-tuning increasingly impractical. Parameter-efficient fine-tuning (PEFT) methods, such as LoRA (lora_original), DoRA (liu2024dora), and HiRA (huang2025hira), address this by injecting trainable low-rank updates into frozen backbones, enabling task adaptation with minimal parameter overhead.
@@
Parameter-efficient fine-tuning (PEFT) methods aim to adapt large-scale pre-trained models to downstream tasks while minimizing the number of trainable parameters. Early approaches include adapter tuning (houlsby2019parameter), prefix-tuning (li2021prefix), and prompt-tuning (lester2021power), which insert task-specific modules or learn prompt embeddings. LoRA (lora_original) further improves efficiency by injecting trainable low-rank matrices into weight projections, achieving strong performance across a range of language and vision tasks. While these methods reduce memory and compute costs, they typically apply uniform updates across all tokens, lacking the flexibility to differentiate token-level importance. In contrast, our method introduces a token-aware gating mechanism that modulates the update strength dynamically per-token, allowing fine-grained control over adaptation.
@@
Table 2 presents the results on the CONVAI2 dialogue generation benchmark. GateRA achieves the best performance on both LLaMA-2-7B and LLaMA-3-8B, with average scores of 47.98 and 48.61, respectively. These results surpass HiRA and all other PEFT baselines across all evaluated metrics, including BLEU, BERT-based F1/Recall/Precision, Meteor, and ROUGE-L. This further confirms the strength of our token-aware gating strategy in controlling adaptation behavior for open-domain generation.
@@
To evaluate the parameter-efficiency of GateRA, we compare its performance under two adaptation ranks: $r=16$ and $r=32$, across eight commonsense reasoning benchmarks. As illustrated in Figure 3, GateRA with $r=16$ achieves an average accuracy of 86.70%, which is highly competitive with the 87.53% obtained by the $r=32$ setting. Remarkably, the lower-rank configuration requires only half the trainable parameters, yet maintains strong task-wise performance across all datasets, with minimal degradation. For instance, GateRA-$r=16$ achieves comparable or even superior performance to the higher-rank baseline on PIQA, ARC-E, and BoolQ. These results highlight the robustness of our token-aware gating mechanism, which effectively prioritizes informative tokens under limited adaptation budgets.

### v3

@@ Introduction
Large pre-trained models have revolutionized natural language processing and vision-language tasks, yet their massive parameter counts make full fine-tuning increasingly impractical. Parameter-efficient fine-tuning (PEFT) methods, such as LoRA (hu2021lora), DoRA (liu2024dora), and HiRA (huang2025hira), address this by injecting trainable low-rank updates into frozen backbones, enabling task adaptation with minimal parameter overhead.
@@
Parameter-efficient fine-tuning (PEFT) methods aim to adapt large-scale pre-trained models to downstream tasks while minimizing the number of trainable parameters. Early approaches include adapter tuning (houlsby2019parameter), prefix-tuning (li2021prefix), and prompt-tuning (lester2021power), which insert task-specific modules or learn prompt embeddings. LoRA (hu2021lora) further improves efficiency by injecting trainable low-rank matrices into weight projections, achieving strong performance across a range of language and vision tasks. While these methods reduce memory and compute costs, they typically apply uniform updates across all tokens, lacking the flexibility to differentiate token-level importance. In contrast, our method introduces a token-aware gating mechanism that modulates the update strength dynamically per-token, allowing fine-grained control over adaptation.
@@
Table 2 presents the results on the CONVAI2 dialogue generation benchmark. GateRA achieves the best performance on both LLaMA-2-7B and LLaMA-3-8B, with average scores of 47.96 and 48.61, respectively. These results surpass HiRA and all other PEFT baselines across all evaluated metrics, including BLEU, BERT-based F1/Recall/Precision, Meteor, and ROUGE-L. This further confirms the strength of our token-aware gating strategy in controlling adaptation behavior for open-domain generation.
@@
To evaluate the parameter-efficiency of GateRA, we compare its performance under two adaptation ranks: $r=16$ and $r=32$, across eight commonsense reasoning benchmarks. As illustrated in Figure 3, GateRA with $r=16$ achieves an average accuracy of 86.70%, which is highly competitive with the 87.53% obtained by the $r=32$ setting. Remarkably, the lower-rank configuration requires only half the trainable parameters, yet maintains strong task-wise performance across all datasets, with minimal degradation. For instance, GateRA with $r=16$ achieves comparable or even superior performance to the higher-rank baseline on PIQA, ARC-E, and BoolQ. These results highlight the robustness of our token-aware gating mechanism, which effectively prioritizes informative tokens under limited adaptation budgets.